can anyone help? -->find the zeros of the equation −3x^2+27x^2+1200=0
Answer by TimothyLamb(4379)

---
-3x^2 + 27x^2 + 1200 = 0
-24x^2 + 1200 = 0
---
the above quadratic equation is in standard form, with a=-24, b=0, and c=1200
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-24 0 1200
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
this quadratic has two real roots (the x-intercepts), which are:
---
answer:
x = -7.07106781
x = 7.07106781
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php